OpenStudy (anonymous):

What is 54.588 in square root form? Please help.

OpenStudy (phi):

do you have an example of what you mean ?

OpenStudy (anonymous):

I'm trying to find the volume of a triangular prism with a height of 7 inches and a base that is an equilateral triangle with sides of 6 inches. My answer so far is 54.588, but 2 of the answers I'm given are in square root form.. I'm just confused. I feel like I've done my work right, but the square roots are throwing me off.

OpenStudy (phi):

It is easiest to start at the beginning. (square roots have decimals that go forever, and 54.588 is an approximation)

OpenStudy (phi):

the area of the equilateral triangle with side = 6 is \[ \frac{s^2 \sqrt{3}}{4} \] using s=6 we get \[ \frac{36 \sqrt{3}}{4} = 9 \sqrt{3}\]

OpenStudy (phi):

now multiply by the height 7 to get the volume \[ 63\sqrt{3} \]

OpenStudy (anonymous):

Then don't you have to divide that by 2? How do you divide a square root by 2?

OpenStudy (phi):

why do you want to divide by 2 ? and yes, you could divide a square root by 2

OpenStudy (phi):

this is a by the way \[\frac{ \sqrt{a}}{b} = \sqrt{\frac{a}{b^2} }\] if you wanted to do that.

OpenStudy (phi):

the are of a triangle is Area = \( \frac{1}{2} b\cdot h\) 1/2 the base times the height but for the equilateral triangle, you can use a different formula \[ Area = \frac{s^2 \sqrt{3}}{4} \]

OpenStudy (anonymous):

Ohh, okay.. so I don't need to divide by 2 at all?

OpenStudy (anonymous):

That's relieving. xD

OpenStudy (anonymous):

Thank you so much. <3